OpenStudy (sushi121212):

Please help me on my system of equations!! w+2x+2y+z=-2 w+3x-2y-z=-6 -2w-x+3y+3z=6 w+4x+y-2z=-14

OpenStudy (sshayer):

w+2x+2y+z=-2 ...(1) w+3x-2y-z=-6 ...(2) -2w-x+3y+3z=6 ...(3) w+4x+y-2z=-14 ...(4) (2)-(1) gives x-4y-2z=-4 ...(5) (3)+2(2) gives 5x-y+z=-6 ...(6) (4)-(2) gives x+3y-z=-8 ...(7) (5)+2(6) gives 11x-6y=-16 ...(8) (6)+(7) gives 6x+2y=-14 ...(9) (8)+3(9) gives 29x=-58 x=-58/29=-2 put x=-2 in (9) -12+2y=-14 2y=-2 y=-1 put x=-2,y=-1 in (6) -10+1+z=-6 z=-6+10-1=3 put x=-2,y=-1,z=3 in (1) w-4-2+3=-2 w=-2+4+2-3=1 x=-2,y=-1,z=3,w=1

OpenStudy (sushi121212):

Oh! Wow!! Thank you for taking the time to do that!!!